A hockey team played 81 games in a season. They won 1 fewer game than three times the number of ties and lost 8 fewer games than they won. How many games did they win? Lose? Tie?: A hockey team played 81 games in a season. They won 1 fewer game than three times the number of ties and lost 8 fewer games than they won. How many games did they win? Lose? Tie?
Answer by josmiceli(2035)

Let number of wins = w
Let number of losses = l
Let number of ties = t
w + l + t = 81
w = 3t - 1
w = l + 8
----------------
3t = w + 1
t = (w+1)/3
I'll get everything in terms of l
l + 8 + l + (w+1)/3 = 81
l + 8 + l + (l+9)/3 = 81
6l + 24 + l + 9 = 243
7l = 243 - 33
7l = 210
l = 30
w = l + 8
w = 38
t = (w+1)/3
t = 39/3
t = 13
There were 38 wins, 30 losses, and 13 ties
check:
w + l + t = 81
38 + 30 + 13 = 81
81 = 81
-------------
w = 3t - 1
38 = 3*13 - 1
38 = 38
-------------
w = l + 8
38 = 30 + 8
38 = 38